OpenStudy (anonymous):

-7-2x^2= -45 Solve quadratic function with square roots

OpenStudy (anonymous):

i am noy sure but ... -7-2x^2=-45 -7-2x^2+45=0 -2x^2+32=0 -2(x^2-16)=0 -2(x^2-16)/2=0/2 x2-16=0 (x+4)(x-4)=0 x=4,-4

OpenStudy (whpalmer4):

Not sure? Try your answer in the original formula, and see if it works! \[-7-2x^2=-45\]we think \(x=4\) might be a solution, so let's try it. \[-7-2(4)^2=-45\]\[-7-2*16 = -45\]\[-7-32=-45\]\[-39=-45\]apparently that's not the solution.

OpenStudy (anonymous):

okey((try you

OpenStudy (anonymous):

so what would the answers be?

OpenStudy (pawanyadav):

-2x^2=-45+7 -2x2=-38 Divided by -2 x2=19 So x=√(19) and -√(19)

OpenStudy (pawanyadav):

Got it...

OpenStudy (anonymous):

yes thank you!

OpenStudy (pawanyadav):

Welcome